OpenStudy (anonymous):

solve the following equation 4 c0s^4 x -1 =0

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

\[4\cos^4(x)-1=0\]

OpenStudy (anonymous):

r u looking for a numerical solution?

OpenStudy (anonymous):

like x= pi + 2npi and 3pi/2 + 2npi

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

Happy Studying!

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

just in case http://www.wolframalpha.com/input/?i=+4+cos%5E4+%28x%29+-1+%3D0

OpenStudy (campbell_st):

why not just do it the old fashion way... before wolfram alpha \[4 \cos^4(x) = 1\] \[\cos^4(x) = \frac{1}{4}\] then \[\cos(x) = \pm \sqrt[4]{\frac{1}{4}}\] so then \[x = \pm \cos^{-1} \frac{1}{\sqrt[4]{4}}\] hope it makes sense

OpenStudy (anonymous):

thank you for your help